4 Derive Snells law from Fermats Principle. Assume that the interface between the two materials is a planer surface
Answer #1

From figure , which shows that the points A and B are in ywo optically different medium with their refractive index n_1 and n_2 respectively. Lets the velocity at the two side of the moundary are v_1 and v_2 respectively. From figure, total time of travel is,

t=\frac{\sqrt{a^2+x^2}}{v_1}+\frac{\sqrt{b^2+(d-x)^2}}{v_2}=\frac{l_1}{v_1}+\frac{l_2}{v_2}

By using relation

t=\frac{n_1l_1+n_2l_2}{c}=\frac{l}{c}

Where l is optical path length. According to fermat principle ,

\frac{dl}{dx}=0

hence

\frac{dl}{dx}=n_1\frac{x}{\sqrt{a^2+x^2}}-n_2\frac{d-x}{\sqrt{b^2+(d-x)^2}}=0\\ \Rightarrow n_1\frac{x}{\sqrt{a^2+x^2}}=n_2\frac{d-x}{\sqrt{b^2+(d-x)^2}}\\ \Rightarrow n_1\sin\theta_1=n_2\sin\theta_2

Which is Snells law.

18 March 201711:23 91 0